Around 1.03 x 10^5 beehives are needed.

Each hive has on average, 5.0 x 10^4 bees.

80% of the beehives are needed.

Find the total of bees needed and write your answer in standard form.


1
Expert's answer
2020-08-31T18:31:05-0400

The total number of bees needed = Required percentage of beehives * total available beehives * average number of bees in a single beehive


The total number of bees needed = "(\frac {80}{100})\ *\ (1.03*10^5)\ *\ (5.0*10^4) = 4.12*10^9"


Hence the total number of bees needed = "4.12*10^9"